Please find 3 consecutive odd integers such that the product of the two smaller integers exceeds the largest integer by 52. (Only an algebraic solution will be excepted)
Your help will be much appreciated

Answer by CubeyThePenguin(3113)

consecutive odd integers: (x-2), x, (x+2)

(x-2)(x) = x+2 + 52
x^2 - 2x = x + 54
x^2 - 3x - 54 = 0
(x - 9)(x + 6) = 0

x is odd, so the integers are 7, 9, and 11.